$A$ student says that if you throw a die,it will show up $1$ or not $1$. Therefore,the probability of getting $1$ and the probability of getting 'not $1$' each is equal to $\frac{1}{2}$. Is this correct? Give reasons.

Vedclass pdf generator app on play store
Vedclass iOS app on app store
(B) No,this is not correct.
When a fair die is thrown,the total number of possible outcomes is $6$,which are ${1, 2, 3, 4, 5, 6}$.
The probability of getting $1$ is $P(1) = \frac{\text{Number of favorable outcomes}}{\text{Total number of outcomes}} = \frac{1}{6}$.
The probability of getting 'not $1$' is $P(\text{not } 1) = 1 - P(1) = 1 - \frac{1}{6} = \frac{5}{6}$.
Since $\frac{1}{6} \neq \frac{5}{6}$,the student's claim that each probability is $\frac{1}{2}$ is incorrect.

Explore More

Similar Questions

$A$ card is drawn from a deck of $52$ cards. The event $E$ is that the card is not an ace of hearts. The number of outcomes favourable to $E$ is

There are $5$ red,$8$ white,and $4$ green balls in a box. $A$ ball is drawn at random from the box. Find the probability that the ball drawn is not green.

If $P(\overline{D}) = \frac{6}{17}$,then $P(D) = \ldots$

There are $5$ white,$7$ red,$4$ black,and $2$ blue identical balls in a box. One ball is selected at random from the box. Find the probability that the selected ball is not white.

It is given that in a group of $3$ students,the probability of $2$ students not having the same birthday is $0.992$. What is the probability that the $2$ students have the same birthday?

Vedclass Products

For Students

Vedclass Test Series

Mock tests in real JEE/NEET style with performance analysis. 5-day free trial.

Start Free Trial
For Teachers

Exam Paper Generator

Generate Set A/B/C/D exam papers from 7.5L+ questions in 2 minutes. 3 chapters free.

Try Free
For Institutes

Online Exam Module

Live online exams with unlimited students, 360° analytics & white-label branding.

See Demo